OpenStudy (anonymous):

-8x+3y=50 -4x-6y=-20 Which of the following is/are solution(s) of the linear system? (-2,-1) (-1,-6) (5,0) (-4,6)

OpenStudy (anonymous):

sorry i am not sure

OpenStudy (zale101):

When you graph those two lines, the intersecting point where it meets on both of the line is the solution of the system.

OpenStudy (zale101):

-8x+3y=50 Is in the standard form of the equation and you’re trying to convert it into the slope-intercept form Ax + By = C standard form Y=mx+b slope intercept form -8x+3y=50 add 8x 3y=8x+50 then divide everything by 3 Y=8/3x+50/3 for line -4x-6y=-20 Is y=(-2/3)x+10/3